# Hybrid Quantum Repeater Chains with Semiconductor Quantum Dots and Group-IV-Vacancy Color Centers in Diamond

### v2

@@ V Discussion and Outlook
For the reported photon bandwidths $\gamma_{\rm X}=4.34$ GHz and $\gamma_{\rm XX}=8.33$ GHz [25], the spin-spin entanglement infidelities within one section are $7.91\times 10^{-2}$ and $6.09\times 10^{-2}$ for the SiV and SnV, respectively. Using spectral filtering we can additionally interface the G4V’s cavity with photons that are narrower than the reported photon bandwidth. For those we find a higher fidelity than without filtering at the same bandwidth due to the steeper shape of the spectrum after the filter compared to a Lorentzian. This shows that temporal photon shaping is an important tool for improving reflective spin-photon interfaces.
@@
Finally, given the versatility of color centers in diamond, we anticipate that our approach is compatible with next-generation repeaters that employ quantum error correction [74, 31, 75].
@@ Appendix A Quantum State Transfer
To demonstrate the principle of quantum state transfer from the photon to a spin [60], we assume an arbitrary photonic state $\alpha|E\rangle+\beta|L\rangle$ and initialize the spin in the $|{1}\rangle$ state. The procedure includes four steps: early-time-bin photon reflection, first $\frac{\pi}{2}$ rotation, late-time-bin photon reflection, and second $\frac{\pi}{2}$ rotation. As a simplified example, we first look at a single-frequency photon, i.e. we assume $|E\rangle=|L\rangle=e^{{\rm i}\omega_{0}t}$ and a half-open cavity without photon loss. The evolution of the combined photon-spin state is as follows.
@@ Appendix B Optimizing Spin–Photon CPHASE Gates
The corresponding amplitude spectrum is lorentzian [57]. We perform the reflection scheme for entangled photons with emission rates $\gamma_{\rm X},\gamma_{\rm XX}$ and spectrum $S_{\rm X},S_{\rm XX}$. It is
@@
with $\alpha,\beta\in\mathbb{C}$ such that $|\alpha|^{2}+|\beta|^{2}=1$. Assuming that every photon gets detected with unity probability the $X$-measurement has the form [76]
@@
using simplicial homology global optimization [77]. In [32] we developed a comprehensive spin-photon interaction model including crosstalk. To optimize subject to crosstalk, we have to convert the integrals shown in Eqs. (37)-(39) to time domain. For that purpose, we define $\mathcal{D}_{k}$ for $k=1,2$ as the map which maps an input signal $a_{\rm in}$ to the reflected signal $a_{\rm out}$ when the spin is initialized in $|k\rangle$. It holds
@@
To optimize the fidelity with crosstalk, we first solve the optimization problem in Eq. (52) without crosstalk. We initialize an optimization method like Nelder-Mead’s algorithm [78] in the respective minimizer found from Eq. (52) and perform the optimization, encountering the time integrals shown in Eqs. (53)-(55).
@@ Appendix C Spin-Spin Entangled State
To model imperfect photon pair generation, we apply a depolarizing channel [79], i.e.
@@
where $\mathcal{D}_{k}(a)$ denotes the output mode $a_{\rm out}$ from the time evolution according to the Heisenberg-Langevin equations shown in [32] initializing $\sigma_{kk}(0)=1$ for an input mode $a_{\rm in}$. After measuring out the photon, the state reads
@@ D.1 Contrast and Cross-Talk
The controlled phase gate between the incident photon and the G4V’s spin is limited by the contrast $\Delta\omega_{s}=\omega_{2B}-\omega_{1A}$. The order of magnitude of the contrast depends on strain and the magnetic field. We choose the magnetic field strengths $B=1$ T and $\theta_{\rm dc}=0$ to analyze the contrast’s dependence on the extrinsic strain of the SiV and SnV. Fig. 6 illustrates the contrast $\Delta\omega_{s}$ as a function of the axial strain $E_{x}$ and shear strain $\epsilon_{xy}$. The dependence of the G4V’s Hamiltonian on strain is shown in [61]. The maximal contrast for the SiV is roughly $\Delta\omega_{s}\approx 1$ GHz at $E_{x}=5.38\cdot 10^{-4}$ and $\epsilon_{xy}=0$. For the SnV, it is $\Delta\omega_{s}\approx 0.8$ GHz for negligible extrinsic strain. From an experimental perspective, it is not possible to control the extrinsic strain in such a precise manner. We observe a small change in the contrast in the local environment of the optimal strain configuration, making the choice of the optimal strain parameters reasonable for the following simulations.
@@
In Fig. 7 we visualize the contrast $\Delta\omega_{s}$, spin splitting $\omega_{s}$, and the cross couplings $g_{1B},g_{2A}$ as a function of the magnetic field orientation $\theta_{dc}$. On the left-hand side, we observe that the contrast decreases for a value of $\theta_{\rm dc}$ until $\theta_{\rm dc}\approx 0.4$. For $\theta_{\rm dc}>0.4$ the contrast increases monotonically until it reaches the value $\Delta\omega_{s}\approx 8(12)$ GHz for the SiV (SnV). The spin splitting decreases monotonically for increasing $\theta_{\rm dc}$ for the SiV and SnV. On the right-hand side, it becomes apparent that the cross-talk reaches its maximum at $\theta_{\rm dc}=\pi/2$ $(1.4)$ for the SiV (SnV). From [32], it becomes apparent that the cross-talk spin splitting ratio is a relevant measure to assess the impact of the cross-talk on the spin-photon interaction. That ratio is
@@ D.2 Optimization
We assume the QD emission rates $\gamma_{X}=4.34$ GHz, $\gamma_{\rm XX}=8.33$ GHz [25] at the optimal strain configuration for the SiV and SnV and the magnetic field strength $B=1$ T. Due to the contrast dependence on the magnetic field orientation, we optimize for spin-photon entanglement for $0\leq\theta_{\rm dc}\leq\pi/2$ for the SiV and SnV to extract an optimal $\theta_{\rm dc}$ for the present configuration.
@@
In Fig. 8 we illustrate the performance of spin-photon entanglement as a function of the magnetic field orientation for the SiV. In Fig. 8a) we illustrate the infidelity curve for $\theta_{\rm dc}\geq 0.7$ because the optical splitting $\Delta\omega_{s}$ is too small to benefit from spin-photon entanglement for $\theta_{\rm dc}<0.7$. For $\theta_{\rm dc}>0.7$, the infidelity decreases, where the infidelity with and without cross-talk are so close that no difference becomes apparent. The decreasing infidelity is due to the rising contrast $\Delta\omega_{s}$. A magnetic field orientation $\theta_{\rm dc}=\pi/2$ produces an optimized infidelity with the value $1-F_{\rm sp}=7.92\cdot 10^{-2}$ including crosstalk. Without crosstalk, the infidelity $1-F_{\rm sp}=8.00\cdot 10^{-2}$ is achieved at the considered magnetic field orientation. That is counterintuitive because crosstalk refers to unwanted couplings, which produce a nonlinear disturbance in the Heisenberg-Langevin equations shown in [32]. However, the optical splitting between the cross-transitions $\omega_{1B}-\omega_{2A}$ is larger than the optical splitting between the resonant transitions $\omega_{1A}-\omega_{2B}$. That means that crosstalk produces a gain of optical splitting in a nonlinear manner, resulting in a potentially lower infidelity. For the present simulations, the crosstalk is small enough that its nonlinear behavior, like frequency mixing, a property of nonlinear oscillations [80], does not deteriorate the infidelity. Local optimization improves the infidelity for the present parameters up to $1-F_{\rm sp}=7.91\cdot 10^{-2}$. In Fig. 8b), the spin-spin entanglement success probability is illustrated as a function of the magnetic field orientation. The efficiencies with and without cross-talk differ so little that they are virtually indistinguishable. The efficiency rises monotonically for $\theta_{\rm dc}>0.7$ since the contrast increases in that range. At the optimal magnetic field orientation, the efficiency is $\eta_{\rm sp}\approx 0.9187$.
@@
In Fig. 9, we illustrate the performance of spin-photon entanglement as a function of the magnetic field orientation for the SnV. The behavior of the infidelity on the left-hand side and efficiency on the right-hand side behaves equally compared to the SiV until the optimal magnetic field orientation $\theta_{\rm dc}=1.285$. For $\theta_{\rm dc}>1.285$, the infidelity with cross-talk rises until $\theta_{\rm dc}\approx 1.4$, where the cross couplings also reach their maximum. The infidelity decreases afterwards because the cross couplings decrease for $\theta_{\rm dc}>1.4$. The efficiency in Fig. 9b) rises monotonically, except near $\theta_{\rm dc}=1.4$. At the orientation $\theta_{\rm dc}=1.33$ we find the infidelity $8.16\cdot 10^{-2}$ with crosstalk and $6.40\cdot 10^{-2}$ without crosstalk. Compared to the SiV crosstalk leads to an increase in the infidelity. However, local optimization using Nelder-Mead’s algorithm [78] is able to improve the infidelity with crosstalk up to $1-F_{\rm sp}=6.09\cdot 10^{-2}$, which is better than the infidelity without crosstalk. That is due to the increase in optical splitting generated by crosstalk. It is not possible to define an optical splitting here because of the nonlinear nature of the Heisenberg-Langevin equations shown in [32]. However, it provides an intuitive understanding of the observed phenomenon. We find the optimal infidelity $1-F_{\rm sp}=6.09\cdot 10^{-2}$ and efficiency $\eta_{\rm sp}=0.9935$.
@@
To provide data which are closer to state-of-the-art cavities and microwave spin control restrictions we assume the magnetic field $B=0.3$ T and $\theta_{\rm dc}=\pi/2$ for the SiV and find $1-F_{\rm sp}<7.37\cdot 10^{-2}$ for $\gamma<1.6$ GHz at $C_{1A}=25$. For that, we assumed crosstalk, which could improve the infidelity from $1-F_{\rm sp}=8.09\cdot 10^{-2}$ to $1-F_{\rm sp}=7.37\cdot 10^{-2}$ using Nelder-Mead’s optimization algorithm [78]. We conclude that broadband QDs are compatible with SiV cavities and magnetic field constraints, enabling high-fidelity spin $\pi/2$ microwave rotations.
For the SnV, we illustrate the spin-spin entanglement infidelity as a function of the cooperativity at $\theta_{\rm dc}=1.33$. In Fig. 10a), we visualize that. We observe that the infidelity decreases for an increasing cooperativity. Due to experimental restrictions [54] we extract the infidelity at $C_{1A},C_{2B}\approx 25$ and find $1-F_{\rm sp}\approx 0.45$. In Fig. 10b), we visualize the infidelity as a function of the cooperativity and bandwidth. We observe that for the fidelity $F_{\rm sp}>0.95$ at $C_{1A}=25$, a bandwidth of $\gamma<0.1$ GHz is required. We conclude that higher cooperativities are of great relevance for the SnV.
@@ D.3 Sensitivity
In Fig. 12 we visualize the optimized amplitude and phase spectra for the SiV and SnV, neglecting cross-talk. For the SnV we use optimized parameters at $\theta_{\rm dc}=\pi/2$ for illustrative purposes. We observe flatness for the phase spectra for the SiV and SnV in a local environment of the optimized central frequency, which is due to robust optimization to broadband photons. We do not illustrate the spectra taking crosstalk into account because of the nonlinear nature of the upcoming terms in the Heisenberg-Langevin equations shown in [32]. Nonlinear systems cannot be described by a single transfer function since a linear input-output relation can only describe LTI (linear time-invariant) systems [81]. It is conceptually possible to illustrate such spectra for a damped input signal on a range of central frequencies. However, they depend on the photon-emitting duration, which is why we omit such visualizations here.
@@ D.4 Filtering
Assuming we use the quantum dot from [25] and aim for narrower photon bandwidths, frequency filtering becomes essential. We model the filter as an empty, symmetric cavity with negligible losses. The transfer function mapping an input mode to a narrower output mode is the transmittivity of an empty cavity [8]
@@
To model spin-photon interaction with crosstalk assuming the photon is emitted from a QD and passing subsequently through a Fabry-Perot interferometer we use the Heisenberg-Langevin equation shown in [32] and the input mode in time domain. To derive the input mode after filtering in time domain $a_{\rm in}$ we evaluate the convolution between the functions
@@
For optimizing a spin-photon CPHASE gate it is sufficient to optimize a cavity which is compatible with the QD from [25]. To evaluate the fidelity and efficiency we assume that the incoming mode is given by Eq. (83) when the model in frequency domain is sufficient and Eq. (86) when time domain simulations are required to capture crosstalk effects and the G4V cavity as well as the emission frequency of the mode do not change. Since $2e_{0}/(2\kappa_{f}-\gamma)$ might be larger than $e_{0}$ one has to carefully choose $e_{0}$ to not produce unwanted driving between the states $|1\rangle,|A\rangle$ and $|2\rangle,|B\rangle$. Furthermore, the one has to multiply the factor $\mathcal{N}_{X,XX}$ from Eqs. (72)-(74) with $\kappa_{f}^{2}$.
@@
Assuming the QD from [25] we apply filtering to make it compatible with the SiV at the magnetic field strength $B=0.3$ T and orientation $\theta_{\rm dc}=\pi/2$. We analyze that case because it leads to a spin splitting $\omega_{s}\approx 8$ GHz which is possible to address with microwaves to produce a spin $\pi/2$ rotation with state-of-the-art experimental techniques [62]. At the field strength $B=1$ T it has not yet been experimentally realized and represents future work in microwave spin control for G4Vs.
@@
To optimize a spin-photon CPHASE gate for the SiV at $B=0.3$ T and $\theta_{\rm dc}=\pi/2$ we first optimize the SiV cavity and central frequency of the emitted photons. For that purpose we are restricted to assume a narrower bandwidth because spin-photon entanglement does not work at $\gamma_{X}=4.34$ GHz and $\gamma_{XX}=8.33$ GHz without filtering at the considered magnetic field strength. We assume a QD with $\gamma_{X}=\gamma_{XX}=1.56$ GHz without filtering to optimize a SiV cavity. By first optimizing the parameters in frequency domain as shown in App. B and subsequently improving the parameters when including crosstalk we find the fidelity $F=7.37\cdot 10^{-2}$ at the parameters $\kappa=17.82$ GHz, $\delta_{c}=\omega_{1A}-\omega_{c}=-38.56$ GHz, $\delta_{0}=\omega_{1A}-\omega_{0}=4.57$ GHz, $C_{1A},C_{2B}\approx 25$. We specifically restricted the cooperativity to $25$ because these are state-of-the-art SiV cavities [54] making the present simulations a meaningful contribution. When assuming the QD from [25] with the rates $\gamma_{X}=4.34$ GHz and $\gamma_{XX}=8.33$ GHz with the above mentioned SiV cavity modeled by $(\kappa,\delta_{c})$ and emission frequency governed by $\delta_{0}$ we find that $\tilde{\gamma}_{X}<4$ GHz and $\tilde{\gamma}_{XX}<7.3$ GHz is sufficient to achieve the infidelities $1-F_{\rm sp}<7.89\cdot 10^{-2}$ and efficiencies $\eta_{\rm sp}<0.3311$. That infidelity bound was used because it is on the same order of magnitude as the optimized spin-spin entanglement infidelity for the SiV at $B=1$ T and the broadband QD from [25] with the rates $\gamma_{X}=4.34$ GHz and $\gamma_{XX}=8.33$ GHz which is shown in Tab. 2. Such a low infidelity is achieved even at so broad spectra using filtering subject to experimental constraints for the magnetic field strength $B=0.3$ T because of the steeper mode spectrum after filtering. That leads to compensation of fidelity loss even if the spectrum is such broad.

### v3

@@ V Discussion and Outlook
For the reported photon bandwidths $\gamma_{\rm X}=4.34$ GHz and $\gamma_{\rm XX}=8.33$ GHz [schimpf_quantum_2021], the spin-spin entanglement infidelities within one section are $7.91\times 10^{-2}$ and $6.09\times 10^{-2}$ for the SiV and SnV, respectively. Using spectral filtering we can additionally interface the G4V’s cavity with photons that are narrower than the reported photon bandwidth. For those we find a higher fidelity than without filtering at the same bandwidth due to the steeper shape of the spectrum after the filter compared to a Lorentzian. This shows that temporal photon shaping is an important tool for improving reflective spin-photon interfaces.
@@
Finally, given the versatility of color centers in diamond, we anticipate that our approach is compatible with next-generation repeaters that employ quantum error correction [borregaard2020one, bopp_sawfish_2024, Ruf2021].
@@ Appendix A Quantum State Transfer
To demonstrate the principle of quantum state transfer from the photon to a spin [borregaard_one-way_2020], we assume an arbitrary photonic state $\alpha|E\rangle+\beta|L\rangle$ and initialize the spin in the $|{1}\rangle$ state. The procedure includes four steps: early-time-bin photon reflection, first $\frac{\pi}{2}$ rotation, late-time-bin photon reflection, and second $\frac{\pi}{2}$ rotation. As a simplified example, we first look at a single-frequency photon, i.e. we assume $|E\rangle=|L\rangle=e^{{\rm i}\omega_{0}t}$ and a half-open cavity without photon loss. The evolution of the combined photon-spin state is as follows.
@@ Appendix B Optimizing Spin–Photon CPHASE Gates
The corresponding amplitude spectrum is lorentzian [tran_nanodiamonds_2017]. We perform the reflection scheme for entangled photons with emission rates $\gamma_{\rm X},\gamma_{\rm XX}$ and spectrum $S_{\rm X},S_{\rm XX}$. It is
@@
with $\alpha,\beta\in\mathbb{C}$ such that $|\alpha|^{2}+|\beta|^{2}=1$. Assuming that every photon gets detected with unity probability the $X$-measurement has the form [propp]
@@
using simplicial homology global optimization [endres_simplicial_2018]. In [strocka_memory_2025] we developed a comprehensive spin-photon interaction model including crosstalk. To optimize subject to crosstalk, we have to convert the integrals shown in Eqs. (37)-(39) to time domain. For that purpose, we define $\mathcal{D}_{k}$ for $k=1,2$ as the map which maps an input signal $a_{\rm in}$ to the reflected signal $a_{\rm out}$ when the spin is initialized in $|k\rangle$. It holds
@@
To optimize the fidelity with crosstalk, we first solve the optimization problem in Eq. (52) without crosstalk. We initialize an optimization method like Nelder-Mead’s algorithm [2020SciPy-NMeth] in the respective minimizer found from Eq. (52) and perform the optimization, encountering the time integrals shown in Eqs. (53)-(55).
@@ Appendix C Spin-Spin Entangled State
To model imperfect photon pair generation, we apply a depolarizing channel [collins_depolarizing-channel_2015], i.e.
@@
where $\mathcal{D}_{k}(a)$ denotes the output mode $a_{\rm out}$ from the time evolution according to the Heisenberg-Langevin equations shown in [strocka_memory_2025] initializing $\sigma_{kk}(0)=1$ for an input mode $a_{\rm in}$. After measuring out the photon, the state reads
@@ D.1 Contrast and Cross-Talk
The controlled phase gate between the incident photon and the G4V’s spin is limited by the contrast $\Delta\omega_{s}=\omega_{2B}-\omega_{1A}$. The order of magnitude of the contrast depends on strain and the magnetic field. We choose the magnetic field strengths $B=1$ T and $\theta_{\rm dc}=0$ to analyze the contrast’s dependence on the extrinsic strain of the SiV and SnV. Fig. 6 illustrates the contrast $\Delta\omega_{s}$ as a function of the axial strain $E_{x}$ and shear strain $\epsilon_{xy}$. The dependence of the G4V’s Hamiltonian on strain is shown in [pieplow_efficient_2024]. The maximal contrast for the SiV is roughly $\Delta\omega_{s}\approx 1$ GHz at $E_{x}=5.38\cdot 10^{-4}$ and $\epsilon_{xy}=0$. For the SnV, it is $\Delta\omega_{s}\approx 0.8$ GHz for negligible extrinsic strain. From an experimental perspective, it is not possible to control the extrinsic strain in such a precise manner. We observe a small change in the contrast in the local environment of the optimal strain configuration, making the choice of the optimal strain parameters reasonable for the following simulations.
@@
In Fig. 7 we visualize the contrast $\Delta\omega_{s}$, spin splitting $\omega_{s}$, and the cross couplings $g_{1B},g_{2A}$ as a function of the magnetic field orientation $\theta_{dc}$. On the left-hand side, we observe that the contrast decreases for a value of $\theta_{\rm dc}$ until $\theta_{\rm dc}\approx 0.4$. For $\theta_{\rm dc}>0.4$ the contrast increases monotonically until it reaches the value $\Delta\omega_{s}\approx 8(12)$ GHz for the SiV (SnV). The spin splitting decreases monotonically for increasing $\theta_{\rm dc}$ for the SiV and SnV. On the right-hand side, it becomes apparent that the cross-talk reaches its maximum at $\theta_{\rm dc}=\pi/2$ $(1.4)$ for the SiV (SnV). From [strocka_memory_2025], it becomes apparent that the cross-talk spin splitting ratio is a relevant measure to assess the impact of the cross-talk on the spin-photon interaction. That ratio is
@@ D.2 Optimization
We assume the QD emission rates $\gamma_{X}=4.34$ GHz, $\gamma_{\rm XX}=8.33$ GHz [schimpf_quantum_2021] at the optimal strain configuration for the SiV and SnV and the magnetic field strength $B=1$ T. Due to the contrast dependence on the magnetic field orientation, we optimize for spin-photon entanglement for $0\leq\theta_{\rm dc}\leq\pi/2$ for the SiV and SnV to extract an optimal $\theta_{\rm dc}$ for the present configuration.
@@
In Fig. 8 we illustrate the performance of spin-photon entanglement as a function of the magnetic field orientation for the SiV. In Fig. 8a) we illustrate the infidelity curve for $\theta_{\rm dc}\geq 0.7$ because the optical splitting $\Delta\omega_{s}$ is too small to benefit from spin-photon entanglement for $\theta_{\rm dc}<0.7$. For $\theta_{\rm dc}>0.7$, the infidelity decreases, where the infidelity with and without cross-talk are so close that no difference becomes apparent. The decreasing infidelity is due to the rising contrast $\Delta\omega_{s}$. A magnetic field orientation $\theta_{\rm dc}=\pi/2$ produces an optimized infidelity with the value $1-F_{\rm sp}=7.92\cdot 10^{-2}$ including crosstalk. Without crosstalk, the infidelity $1-F_{\rm sp}=8.00\cdot 10^{-2}$ is achieved at the considered magnetic field orientation. That is counterintuitive because crosstalk refers to unwanted couplings, which produce a nonlinear disturbance in the Heisenberg-Langevin equations shown in [strocka_memory_2025]. However, the optical splitting between the cross-transitions $\omega_{1B}-\omega_{2A}$ is larger than the optical splitting between the resonant transitions $\omega_{1A}-\omega_{2B}$. That means that crosstalk produces a gain of optical splitting in a nonlinear manner, resulting in a potentially lower infidelity. For the present simulations, the crosstalk is small enough that its nonlinear behavior, like frequency mixing, a property of nonlinear oscillations [Vanhille2014], does not deteriorate the infidelity. Local optimization improves the infidelity for the present parameters up to $1-F_{\rm sp}=7.91\cdot 10^{-2}$. In Fig. 8b), the spin-spin entanglement success probability is illustrated as a function of the magnetic field orientation. The efficiencies with and without cross-talk differ so little that they are virtually indistinguishable. The efficiency rises monotonically for $\theta_{\rm dc}>0.7$ since the contrast increases in that range. At the optimal magnetic field orientation, the efficiency is $\eta_{\rm sp}\approx 0.9187$.
@@
In Fig. 9, we illustrate the performance of spin-photon entanglement as a function of the magnetic field orientation for the SnV. The behavior of the infidelity on the left-hand side and efficiency on the right-hand side behaves equally compared to the SiV until the optimal magnetic field orientation $\theta_{\rm dc}=1.285$. For $\theta_{\rm dc}>1.285$, the infidelity with cross-talk rises until $\theta_{\rm dc}\approx 1.4$, where the cross couplings also reach their maximum. The infidelity decreases afterwards because the cross couplings decrease for $\theta_{\rm dc}>1.4$. The efficiency in Fig. 9b) rises monotonically, except near $\theta_{\rm dc}=1.4$. At the orientation $\theta_{\rm dc}=1.33$ we find the infidelity $8.16\cdot 10^{-2}$ with crosstalk and $6.40\cdot 10^{-2}$ without crosstalk. Compared to the SiV crosstalk leads to an increase in the infidelity. However, local optimization using Nelder-Mead’s algorithm [2020SciPy-NMeth] is able to improve the infidelity with crosstalk up to $1-F_{\rm sp}=6.09\cdot 10^{-2}$, which is better than the infidelity without crosstalk. That is due to the increase in optical splitting generated by crosstalk. It is not possible to define an optical splitting here because of the nonlinear nature of the Heisenberg-Langevin equations shown in [strocka_memory_2025]. However, it provides an intuitive understanding of the observed phenomenon. We find the optimal infidelity $1-F_{\rm sp}=6.09\cdot 10^{-2}$ and efficiency $\eta_{\rm sp}=0.9935$.
@@
To provide data which are closer to state-of-the-art cavities and microwave spin control restrictions we assume the magnetic field $B=0.3$ T and $\theta_{\rm dc}=\pi/2$ for the SiV and find $1-F_{\rm sp}<7.37\cdot 10^{-2}$ for $\gamma<1.6$ GHz at $C_{1A}=25$. For that, we assumed crosstalk, which could improve the infidelity from $1-F_{\rm sp}=8.09\cdot 10^{-2}$ to $1-F_{\rm sp}=7.37\cdot 10^{-2}$ using Nelder-Mead’s optimization algorithm [2020SciPy-NMeth]. We conclude that broadband QDs are compatible with SiV cavities and magnetic field constraints, enabling high-fidelity spin $\pi/2$ microwave rotations.
For the SnV, we illustrate the spin-spin entanglement infidelity as a function of the cooperativity at $\theta_{\rm dc}=1.33$. In Fig. 10a), we visualize that. We observe that the infidelity decreases for an increasing cooperativity. Due to experimental restrictions [bhaskar_experimental_2020] we extract the infidelity at $C_{1A},C_{2B}\approx 25$ and find $1-F_{\rm sp}\approx 0.45$. In Fig. 10b), we visualize the infidelity as a function of the cooperativity and bandwidth. We observe that for the fidelity $F_{\rm sp}>0.95$ at $C_{1A}=25$, a bandwidth of $\gamma<0.1$ GHz is required. We conclude that higher cooperativities are of great relevance for the SnV.
@@ D.3 Sensitivity
In Fig. 12 we visualize the optimized amplitude and phase spectra for the SiV and SnV, neglecting cross-talk. For the SnV we use optimized parameters at $\theta_{\rm dc}=\pi/2$ for illustrative purposes. We observe flatness for the phase spectra for the SiV and SnV in a local environment of the optimized central frequency, which is due to robust optimization to broadband photons. We do not illustrate the spectra taking crosstalk into account because of the nonlinear nature of the upcoming terms in the Heisenberg-Langevin equations shown in [strocka_memory_2025]. Nonlinear systems cannot be described by a single transfer function since a linear input-output relation can only describe LTI (linear time-invariant) systems [Trentelman2002]. It is conceptually possible to illustrate such spectra for a damped input signal on a range of central frequencies. However, they depend on the photon-emitting duration, which is why we omit such visualizations here.
@@ D.4 Filtering
Assuming we use the quantum dot from [schimpf_quantum_2021] and aim for narrower photon bandwidths, frequency filtering becomes essential. We model the filter as an empty, symmetric cavity with negligible losses. The transfer function mapping an input mode to a narrower output mode is the transmittivity of an empty cavity [reiserer_cavity-based_2015]
@@
To model spin-photon interaction with crosstalk assuming the photon is emitted from a QD and passing subsequently through a Fabry-Perot interferometer we use the Heisenberg-Langevin equation shown in [strocka_memory_2025] and the input mode in time domain. To derive the input mode after filtering in time domain $a_{\rm in}$ we evaluate the convolution between the functions
@@
For optimizing a spin-photon CPHASE gate it is sufficient to optimize a cavity which is compatible with the QD from [schimpf_quantum_2021]. To evaluate the fidelity and efficiency we assume that the incoming mode is given by Eq. (83) when the model in frequency domain is sufficient and Eq. (86) when time domain simulations are required to capture crosstalk effects and the G4V cavity as well as the emission frequency of the mode do not change. Since $2e_{0}/(2\kappa_{f}-\gamma)$ might be larger than $e_{0}$ one has to carefully choose $e_{0}$ to not produce unwanted driving between the states $|1\rangle,|A\rangle$ and $|2\rangle,|B\rangle$. Furthermore, the one has to multiply the factor $\mathcal{N}_{X,XX}$ from Eqs. (72)-(74) with $\kappa_{f}^{2}$.
@@
Assuming the QD from [schimpf_quantum_2021] we apply filtering to make it compatible with the SiV at the magnetic field strength $B=0.3$ T and orientation $\theta_{\rm dc}=\pi/2$. We analyze that case because it leads to a spin splitting $\omega_{s}\approx 8$ GHz which is possible to address with microwaves to produce a spin $\pi/2$ rotation with state-of-the-art experimental techniques [karapatzakis_microwave_2024]. At the field strength $B=1$ T it has not yet been experimentally realized and represents future work in microwave spin control for G4Vs.
@@
To optimize a spin-photon CPHASE gate for the SiV at $B=0.3$ T and $\theta_{\rm dc}=\pi/2$ we first optimize the SiV cavity and central frequency of the emitted photons. For that purpose we are restricted to assume a narrower bandwidth because spin-photon entanglement does not work at $\gamma_{X}=4.34$ GHz and $\gamma_{XX}=8.33$ GHz without filtering at the considered magnetic field strength. We assume a QD with $\gamma_{X}=\gamma_{XX}=1.56$ GHz without filtering to optimize a SiV cavity. By first optimizing the parameters in frequency domain as shown in App. B and subsequently improving the parameters when including crosstalk we find the fidelity $F=7.37\cdot 10^{-2}$ at the parameters $\kappa=17.82$ GHz, $\delta_{c}=\omega_{1A}-\omega_{c}=-38.56$ GHz, $\delta_{0}=\omega_{1A}-\omega_{0}=4.57$ GHz, $C_{1A},C_{2B}\approx 25$. We specifically restricted the cooperativity to $25$ because these are state-of-the-art SiV cavities [bhaskar_experimental_2020] making the present simulations a meaningful contribution. When assuming the QD from [schimpf_quantum_2021] with the rates $\gamma_{X}=4.34$ GHz and $\gamma_{XX}=8.33$ GHz with the above mentioned SiV cavity modeled by $(\kappa,\delta_{c})$ and emission frequency governed by $\delta_{0}$ we find that $\tilde{\gamma}_{X}<4$ GHz and $\tilde{\gamma}_{XX}<7.3$ GHz is sufficient to achieve the infidelities $1-F_{\rm sp}<7.89\cdot 10^{-2}$ and efficiencies $\eta_{\rm sp}<0.3311$. That infidelity bound was used because it is on the same order of magnitude as the optimized spin-spin entanglement infidelity for the SiV at $B=1$ T and the broadband QD from [schimpf_quantum_2021] with the rates $\gamma_{X}=4.34$ GHz and $\gamma_{XX}=8.33$ GHz which is shown in Tab. 2. Such a low infidelity is achieved even at so broad spectra using filtering subject to experimental constraints for the magnetic field strength $B=0.3$ T because of the steeper mode spectrum after filtering. That leads to compensation of fidelity loss even if the spectrum is such broad.